OpenStudy (anonymous):

x^(2)- 3 / x^(2)-6x+5 = 2x+3 / x-5 - x+3 / x-1 = o

OpenStudy (anonymous):

there is no solution that exists

OpenStudy (anonymous):

is this \[\frac{x^2-3}{x^2-6x+5}=\frac{2x+3}{x-5}-\frac{x+3}{x-1}=0\]? thre are too many minus signs here